HRMS, 28th Colloquium, Dijon, France, Aug. 28 - Sept. 1, 2023

ROTATIONAL SPECTROSCOPY OF MOLECULAR IONS USING A NOVEL DOUBLE RESONANCE APPROACH

\underline{O. ASVANY}, P. C. SCHMID, W. SILVA, D. GUPTA, T. SALOMON, S. THORWIRTH, S. SCHLEMMER, I. Physikalisches Institut, Universität zu Köln,

Zülpicher Str. 77, 50937 Köln, Germany

A novel method of rotational spectroscopy of trapped, cold molecular ions is presented. It is a double resonance scheme combining rotational excitation (in the mmwave region) with rovibrational excitation (in the IR). The rovibrational excitation is detected by the recently developed leak-out spectroscopy (LOS) approach^{1}^{,}^{2}. LOS is based on the vibrational-translational energy transfer process occuring during the collision of the IR excited ion with a noble gas atom, after which the ion may escape the trap (”leak out”) so that it can be counted by a detector. LOS and the rotational scheme based on it are very sensitive and generally applicable to any cation or anion. This poster presents rotational spectra for HC_{3}O^{+}, c-C_{3}H_{2}D^{+} and H_{2}CCCH^{+}.

1doi:10.1021/acs.jpca.2c05767, P. C. Schmid, O. Asvany, T. Salomon, S. Thorwirth, and S. Schlemmer, J. Phys. Chem. A, 126, 8111-8117, (2022).

2see also talk of P. C. Schmid in Minisymposium ”Spectroscopy of ions in traps”

p-number: p001  Submitted on Fri, 21 Apr 2023 18:54:20 +0200